# Lie symmetry analysis of (2+1)-dimensional time fractional Kadomtsev-Petviashvili equation

### v2

@@
Abstract: In this paper, Lie symmetry analysis method is applied to the (2+1)-dimensional time fractional Kadomtsev-Petviashvili (KP) equation with the mixed derivative of Riemann-Liouville time-fractional derivative and integer-order $x$-derivative. We obtained all the Lie symmetries admitted by the KP equation and used them to reduce the (2+1)-dimensional fractional partial differential equation with Riemann-Liouville fractional derivative to some (1+1)-dimensional fractional partial differential equations with Erdélyi-Kober fractional derivative or Riemann-Liouville fractional derivative, thereby getting some exact solutions of the reduced equations. In addition, the new conservation theorem and the generalization of Noether operators are developed to construct the conservation laws for the equation studied.
Keywords: Lie symmetry analysis; fractional Kadomtsev-Petviashvili equation; Riemann-Liouville fractional derivative; Erdélyi-Kober fractional derivative; conservation laws.
Mathematics Subject Classification 2020: 76M60; 35G50; 37C79; 34K37.
1. Introduction
@@
which is one of the few integrable equations in high dimensions and comes from the study of long gravity waves in a single layer or multilayered shallow fluid when the waves propagate predominantly in one direction with a small perturbation in the perpendicular one [2]. It also appears in many other fields, such as plasma physics, gas dynamics, etc. Due to the widespread application of Eq. (1.1), it has attracted the interest of many scholars. Especially, Chen [3] found the Backlund transformations related to symmetries and integrability of the KP equation in 1975. Cheng and Li [4] discussed the constraint of the KP equation and its special solutions. Lou [5] derived an explicit and simple constructive formula for the symmetries of the KP equation. Biondini and Pelinovsky [6] discussed the two-dimensional localized solution and resonant two-soliton solution and their properties for the KP equation. Ma [7] presented a class of lump solutions, rationally localized in all directions in the space, to the KP equation through symbolic computation with Maple. In addition, a large number of research results on the KP equation have emerged in recent years (see [8–15] and the references therein).
@@
for $t>a$. We denote the operator ${}_{0}D^{\alpha}_{t}$ as $D^{\alpha}_{t}$ throughout this paper, while $D^{-\alpha}_{t}=I^{\alpha}_{t}$ is Riemann-Liouville fractional integral. Recently, many researchers have used different methods and techniques to study the fractional KP equation (see [17–24] and references therein). Especially, Borluk et al. [22] proved the existence of periodically modulated solitary wave solutions of the fractional KP equation by dimension-breaking bifurcation and discussed the line solitary wave solutions and their transverse (in)stability. They also proved existence of lump solutions for the fractional KP equation via a variational approach, and presented numerically generated lump solutions and observed the cross-sectional symmetry of the solutions numerically in [23].
As a generalization of the classical calculus, fractional calculus can be traced back to the letter written by L’Hôspital to Leibniz in 1695. Since then, it has gradually gained the attention of mathematicians. Especially in recent decades, it has developed rapidly and been successfully applied in many fields of science and technology [25–28]. Therefore, it is very important to find the solution of fractional differential equation. So far, there have been some numerical and analytical methods, such as Adomian decomposition method [29], finite difference method [30], homotopy perturbation method [31], the sub-equation method [32], the variational iteration method [33], Lie symmetry analysis method [34], invariant subspace method [35] and so on. Among them, Lie symmetry analysis method has received an increasing attention.
Lie symmetry analysis method was founded by Norwegian mathematician Sophus Lie at the end of the nineteenth century and then further developed by some other mathematicians, such as Ovsiannikov [36], Olver [37], Ibragimov [38–40] and so on. As a modern method among many analytic techniques, Lie symmetry analysis has been extended to fractional differential equations (FDEs) by Gazizov et al. [34] in 2007. It was then effectively applied to various models of the FDEs occurring in different areas of applied science (see [41–50]). Lie symmetry analysis method can treat differential equations uniformly regardless of their forms, transforming some solutions of these equations into other forms of solutions by means of continuous point transformations [51]. The algorithm of its using for fractional partial differential equations (FPDEs) is as follows:
@@
3. Similarity reductions and invariant solutions of Eq. (1.2)
@@
where $10<\alpha<$ 1 and $f(t)$ is Lipschitz continuous on $[a,b]$. Coincidentally, the integrals in the sense of the Hadamard’s finite-part integral has the same results as the case of convergence [52–54].
@@
Firstly, from the fundamental operator identity, i.e., (see [57–58])
@@ 5. Conclusion
Confict of Interest
The authors have no competing interests to declare that are relevant to the content of this article.

### v3

# Open Communications in Nonlinear Mathematical Physics   ]ocnmp[  Vol.4 (2024) pp 1References   Article

@@
Abstract
In this paper, Lie symmetry analysis method is applied to the (2+1)-dimensional time fractional Kadomtsev-Petviashvili (KP) equation with the mixed derivative of Riemann-Liouville time-fractional derivative and integer-order $x$-derivative. We obtained all the Lie symmetries admitted by the KP equation and used them to reduce the (2+1)-dimensional fractional partial differential equation with Riemann-Liouville fractional derivative to some (1+1)-dimensional fractional partial differential equations with Erdélyi-Kober fractional derivative or Riemann-Liouville fractional derivative, thereby getting some exact solutions of the reduced equations. In addition, the new conservation theorem and the generalization of Noether operators are developed to construct the conservation laws for the equation studied.
1 Introduction
@@
which is one of the few integrable equations in high dimensions and comes from the study of long gravity waves in a single layer or multilayered shallow fluid when the waves propagate predominantly in one direction with a small perturbation in the perpendicular one [2]. It also appears in many other fields, such as plasma physics, gas dynamics, etc. Due to the widespread application of Eq. (1.1), it has attracted the interest of many scholars. Especially, Chen [3] found the Backlund transformations related to symmetries and integrability of the KP equation in 1975. Cheng and Li [4] discussed the constraint of the KP equation and its special solutions. Lou [5] derived an explicit and simple constructive formula for the symmetries of the KP equation. Biondini and Pelinovsky [6] discussed the two-dimensional localized solution and resonant two-soliton solution and their properties for the KP equation. Ma [7] presented a class of lump solutions, rationally localized in all directions in the space, to the KP equation through symbolic computation with Maple. In addition, a large number of research results on the KP equation have emerged in recent years (see [8, 9, 10, 11, 12, 13, 14, 15] and the references therein).
@@
for $t>a$. We denote the operator ${}_{0}D^{\alpha}_{t}$ as $D^{\alpha}_{t}$ throughout this paper, while $D^{-\alpha}_{t}=I^{\alpha}_{t}$ is Riemann-Liouville fractional integral. Recently, many researchers have used different methods and techniques to study the fractional KP equation (see [17, 18, 19, 20, 21, 22, 23, 24] and references therein). Especially, Borluk et al. [22] proved the existence of periodically modulated solitary wave solutions of the fractional KP equation by dimension-breaking bifurcation and discussed the line solitary wave solutions and their transverse (in)stability. They also proved existence of lump solutions for the fractional KP equation via a variational approach, and presented numerically generated lump solutions and observed the cross-sectional symmetry of the solutions numerically in [23].
As a generalization of the classical calculus, fractional calculus can be traced back to the letter written by L’Hôspital to Leibniz in 1695. Since then, it has gradually gained the attention of mathematicians. Especially in recent decades, it has developed rapidly and been successfully applied in many fields of science and technology [25, 26, 27, 28]. Therefore, it is very important to find the solution of fractional differential equation. So far, there have been some numerical and analytical methods, such as Adomian decomposition method [29], finite difference method [30], homotopy perturbation method [31], the sub-equation method [32], the variational iteration method [33], Lie symmetry analysis method [34], invariant subspace method [35] and so on. Among them, Lie symmetry analysis method has received an increasing attention.
Lie symmetry analysis method was founded by Norwegian mathematician Sophus Lie at the end of the nineteenth century and then further developed by some other mathematicians, such as Ovsiannikov [36], Olver [37], Ibragimov [38, 39, 40] and so on. As a modern method among many analytic techniques, Lie symmetry analysis has been extended to fractional differential equations (FDEs) by Gazizov et al. [34] in 2007. It was then effectively applied to various models of the FDEs occurring in different areas of applied science (see [41, 42, 43, 44, 45, 46, 47, 48, 49, 50]). Lie symmetry analysis method can treat differential equations uniformly regardless of their forms, transforming some solutions of these equations into other forms of solutions by means of continuous point transformations [51]. The algorithm of its using for fractional partial differential equations (FPDEs) is as follows:
@@
3 Similarity reductions and invariant solutions of Eq. (1.2)
@@
where $10<\alpha<$ 1 and $f(t)$ is Lipschitz continuous on $[a,b]$. Coincidentally, the integrals in the sense of the Hadamard’s finite-part integral has the same results as the case of convergence [52, 53, 54].
@@
Firstly, from the fundamental operator identity, i.e., (see [57, 58])
@@ 5 Conclusion
Acknowledgements
The authors would like to thank the anonymous reviewers for their careful reading of the previous version of the paper and for their comments and suggestions which have improved the paper very much.